# CEHR-GPT: Generating Electronic Health Records with Chronological Patient Timelines

### v1

@@ Abstract
Synthetic Electronic Health Records (EHR) have emerged as a pivotal tool in advancing healthcare applications and machine learning models, particularly for researchers without direct access to healthcare data. Although existing methods, like rule-based approaches and generative adversarial networks (GANs), generate synthetic data that resembles real-world EHR data, these methods often use a tabular format, disregarding temporal dependencies in patient histories and limiting data replication. Recently, there has been a growing interest in leveraging Generative Pre-trained Transformers (GPT) for EHR data. This enables applications like disease progression analysis, population estimation, counterfactual reasoning, and synthetic data generation. In this work, we focus on synthetic data generation and demonstrate the capability of training a GPT model using a particular patient representation derived from CEHR-BERT, enabling us to generate patient sequences that can be seamlessly converted to the Observational Medical Outcomes Partnership (OMOP) data format.
@@ 1 Introduction
Synthetic data is not real data, that is, it doesn’t relate to any specific individual. However, it mimics the statistical characteristics and journeys of specific patient populations. Synthetic data enables a broader range of researchers to answer their questions of interest without going through the cumbersome process of accessing real data and worrying about patient privacy [2]. In recent years, many machine learning, specifically deep learning and generative artificial intelligence (AI) models, have been developed to derive synthetic data from real EHR data. [3] However, most existing methods for synthetic EHR data generation fail to adequately capture the temporal dependencies that are often critical in medical scenarios. These temporal aspects, such as medication schedules, symptom progression, and lab result timelines, are vital for a comprehensive understanding of patient health trajectories and for developing effective treatment strategies. An ideal synthetic dataset derived from institutional data should maintain the inherent correlations among time-series features, thus enabling researchers to externally validate machine learning models in different populations. Crucially, the synthetic dataset must preserve accurate patient timelines, as predictive tasks are highly susceptible to temporal variations. A synthetic dataset is considered to exhibit comparable machine learning utility to the original data if it meets two key criteria: 1) it demonstrates similar outcome prevalence to the source data; 2) machine learning models trained on the synthetic data achieve performance metrics akin to those trained with the original data.
@@
  • We design a novel patient representation that not only keeps track of the visit types and discharge facility for inpatient visits but also preserves all temporal information including patients’ starting year, starting age, time intervals between visits, and inpatient visit spans. To the best of our knowledge, this is the first time that temporal information is fully retained.
@@ 2 Related work
Until now, almost all the existing approaches used some variation of GAN for learning the data distribution, but unfortunately, GANs are notoriously difficult to train and easily subject to mode collapse. Despite the recent advancements in optimization techniques such as Wasserstein-GAN [12], it would require a significant amount of time to tune hyperparameters to properly train GANs. As the previous works have demonstrated [13, 14, 15], patient sequence generation could be conceptually represented as a language modeling problem. Foresight [16] adapted GPT to forecast patient trajectories. Their methodused a name entity recognition (NER) tool to extract medical concepts from discharge summaries, based on which a patient sequence was constructed chronologically, and then they trained a standard GPT model using all patient sequences constructed from the previous step. To forecast future events, they fed a patient history to prompt the model and employed a Monte Carlo sampling strategy to calculate the probability of developing certain conditions. One limitation of this work was that the model could not predict when a certain condition would happen due to a lack of temporal information in their patient sequence. Nevertheless, their work demonstrated the potential of using GPT for modeling patient sequences. We aim to address this limitations in our work.
@@ 3 Methods
In Figure 1, we present a framework for generating synthetic EHR data from an OMOP source. To retain the temporal dependencies, we opted to work directly with time-series patient sequences instead of using a bag-of-words (BOW) representation by building on our previous work [13]. We first encoded the OMOP data into patient sequences using a specific patient representation, described later in this section. Secondly, we trained a generative model on the converted patient sequences and utilized it to generate new synthetic patient sequences.
[Uncaptioned image] Figure 1: Overall architecture. The OMOP data is first converted to patient sequences by an OMOP encoder based on the patient representation that preserves demographics, visit types, and temporal intervals between visits. Then a generative model is trained to learn the sequence distribution in order to generate new sequences. Next, the generated sequences are converted back to the OMOP format using an OMOP decoder.
Finally, we fed the synthetic patient sequences into an OMOP decoder to create a synthetic OMOP dataset. Furthermore, an evaluation procedure was developed to assess the similarity between the synthetic OMOP and the source OMOP data.
@@ 4.2.1 Dimension-wise Distribution
[Uncaptioned image] Figure 3: KL divergence for comparing concept probability distribution between synthetic data and real data. The probabilities of concepts were calculated on the scale of the entire population.
@@
[Uncaptioned image] Figure 4: Concept prevalence comparison between the source OMOP and generated OMOP using top $p=95\%$ in the log scale stratified by domain in columns and by population in rows, where x-axis and y-axis represent the source and the synthetic data respectively, and each dot represents a concept
@@ 4.2.2 Co-occurrence Relationship
[Uncaptioned image] Figure 5: KL divergence associated with different synthetic data. The closer to the lower bound in the bottom left corner, the better the synthetic data.
[Uncaptioned image] Figure 6: Top 100 co-occurring concept pairs for each co-occurrence category e.g. condition-condition interpreted as a condition concept followed by another condition concept. The x and y axes represent the synthetic and source data respectively
@@
[…]
@@ 4.2.3 Predictive performance
HF readmission
HF patients who have a 30-day all-cause readmission. Observation window: 360 days, Prediction windows 30 days
Hospitalization
2-year risk of hospitalization starting from the 3rd year since the initial entry into the EHR system. Observation window: 540 days, hold-off window: 180 days, prediction windows: 720 days
COPD readmission
COPD patients who have a 30-day all-cause readmission. Observation window: 360 days, prediction windows: 30 days
Afib ischemic stroke
Afib patients with 1-year risk since the initial diagnosis of afib ischemic stroke. Observation window: 720 days, prediction windows: 360 days
CAD CABG
Patients initially diagnosed with Coronary Arterial Disease (CAD) without any prior stent graft will receive the Coronary artery bypass surgery (CABG) treatment. Observation window: 720 days, prediction windows: 360 days
@@
[Uncaptioned image] Figure 8: The consolidated distance metrics for different synthetic datasets stratified by cohort. GPT-Vanilla and CEHR-BERT were omitted as the cohorts couldn’t be constructed due to the loss of temporal information. GPT-OUTPAT was omitted from HF readmission and COPD readmission as the cohorts for the same reason.
@@ 4.3 Privacy Evaluations
We adopted the privacy evaluation framework outlined in [17], focusing on quantifying the risk of privacy breaches through attribute and member inference attacks. The attribute inference risk assesses the likelihood of inferring sensitive attributes from a synthetic dataset by utilizing targeted patients’ demographic information and common diagnoses found in real datasets available to the adversary. In a membership inference attack, the adversary tries to determine if the target record was in the training set given all or partial attributes.
@@ 4.3.1 Membership Inference Attack
We simulated two different attack scenarios, namely the dataset attack and the model attack. In a dataset attack, attackers only have access to the synthetic dataset, whereas they can query the model itself in a model attack. Their goal is to find a way to infer whether or not a real patient record was used for training [19]. For example, the model might output a high probability associated with the training data points and a low probability for the non-training ones, similarly, the synthetic data might be more similar to the training data than the non-training ones.
The attack dataset was constructed based on the following procedure 1) we created the negative examples by taking the 75000 holdout set and labeling them as negative; 2) we created the positive examples by randomly selecting 75000 patients from the training set and assigning them as positive. In the dataset attack, for each attack data point, we found the best match from the synthetic data using a hamming distance metric, which sums up the absolute differences in age, number of visits, and concepts [19], finally, the median distance was used as a threshold, below which patients were predicted as positive and negative otherwise. In the model attack, we computed the average loss of each patient sequence by feeding the attack dataset to the model directly, similarly, the median loss was used to infer the positive/negative cases, where the positive cases were assigned to those below the threshold and positive otherwise. Finally, we calculated recall/precision/f1/accuracy using the ground truth labels against those inferred predictions. Table 4 shows that the accuracy of both attacks is slightly less than 50%, indicating that the performance of such attacks is worse than a random guess.
Accuracy Recall Precision F1
Model Attack 0.4955 0.4955 0.4942 0.4949
Data Attack 0.4941 0.4996 0.4959 0.4978
Table 4: Membership Inference Attack metrics
@@ 4.3.2 Attribute Inference Attack
In attribute inference attack, we assumed the attacker has access to a group of target patients along with their demographic data, including age, gender, race, the year of their initial clinical visit, and prevalent clinical conditions like hypertension, abdominal pain, chest pain, etc. With these features, the attacker identifies the patient in the synthetic dataset with the closest attribute resemblance to the target individual. Then the attacker aims to use the sensitive attributes of matched patient from synthetic dataset to infer the corresponding sensitive attributes of the target patient.
@@
To quantitatively evaluate the attribute inference risk, we randomly sampled 150,000 patients from the real patient dataset as a target group. The search group compromised 1 million generated synthetic patients. To find most similar patients from the search group, we created a set of common attributes including the demographic data along with the top 1% of the most prevalent condition concepts, represented as one-hot encoded features. Then a k-nearest neighbors (KNN with k = 1) algorithm was applied to each target patient and a synthetic patient with the smallest euclidean distance was found. Finally we extracted the sensitive attributes (condition concepts not in the top 1% tier) from matched target and synthetic patients. F1 scores were computed for the sensitive attributes of each matched patient pair and aggregated across all matched patients.
@@
A baseline analysis was performed substituting real patients for synthetic patients. A result lower than the baseline suggests that the likelihood of finding a synthetic data similar to a real patient is lower than finding a real patient similar to a real patient who share sensitive attributes. This implies a lower attribute inference risk which could be acceptable. Our training set was randomly divided into two halves, with 1 million real patients assigned as the target group, and the remaining 1 million as the search group. The matching process was consistently applied and an aggregated F1 score was computed. Table 5 shows that the F1 score of synthetic vs real is less than real vs real scenario. But it’s still higher than other models who didn’t capture the real patterns as effectively. [19]
Recall Precision F1
Synthetic vs Real 0.0350 0.0343 0.0271
Real vs Real 0.0612 0.0468 0.0421
Table 5: Attribute Inference Attack metrics

### v2

@@ Abstract
Synthetic Electronic Health Records (EHR) have emerged as a pivotal tool in advancing healthcare applications and machine learning models, particularly for researchers without direct access to healthcare data. Although existing methods, like rule-based approaches and generative adversarial networks (GANs), generate synthetic data that resembles real-world EHR data, these methods often use a tabular format, disregarding temporal dependencies in patient histories and limiting data replication. Recently, there has been a growing interest in leveraging Generative Pre-trained Transformers (GPT) for EHR data. This enables applications like disease progression analysis, population estimation, counterfactual reasoning, and synthetic data generation. In this work, we focus on synthetic data generation and demonstrate the capability of training a GPT model using a particular patient representation derived from CEHR-BERT, enabling us to generate patient sequences, which can be seamlessly converted to the Observational Medical Outcomes Partnership (OMOP) data format.
@@ 1 Introduction
Synthetic data is not real data, that is, it doesn’t relate to any specific individual. However, it mimics the statistical characteristics and journeys of specific patient populations. Synthetic data enables a broader range of researchers to answer their questions of interest without going through the cumbersome process of accessing real data and worrying about patient privacy [2]. In recent years, many machine learning, specifically deep learning and generative artificial intelligence (AI) models, have been developed to derive synthetic data from real EHR data [3]. However, most existing methods for synthetic EHR data generation fail to adequately capture the temporal dependencies that are often critical in medical scenarios. These temporal aspects, such as medication schedules, symptom progression, and lab result timelines, are vital for a comprehensive understanding of patient health trajectories and for developing effective treatment strategies. An ideal synthetic dataset derived from institutional data should maintain the inherent correlations among time-series features, thus enabling researchers to externally validate machine learning models in different populations. Crucially, the synthetic dataset must preserve accurate patient timelines, as predictive tasks are highly susceptible to temporal variations. A synthetic dataset is considered to exhibit comparable machine learning utility to the original data if it meets two key criteria: 1) it demonstrates similar outcome prevalence to the source data; 2) machine learning models trained on the synthetic data achieve performance metrics akin to those trained with the original data.
@@
  • We design a novel patient representation that captures visit types, discharge facilities for inpatient visits, and all temporal data, such as starting year, age, intervals between visits, and inpatient visit duration. This is the first instance of fully preserving such temporal information, to our knowledge.
@@
  • We utilized the state-of-the-art evaluation framework proposed by Yan et al. [10] to quantify the privacy risk by calculating a number of privacy metrics for synthetic Electronic Health Records (EHR) data, and demonstrated that the privacy risk associated with this synthetic data is low.
@@ 2 Related work
Until now, almost all the existing approaches used some variation of GAN for learning the data distribution, but unfortunately, GANs are notoriously difficult to train and easily subject to mode collapse. Despite the recent advancements in optimization techniques such as Wasserstein-GAN [13], it would require a significant amount of time to tune hyperparameters to train GANs. As the previous works have demonstrated [14, 15, 16], patient sequence generation could be conceptually represented as a language modeling problem. Foresight [17] adapted GPT to forecast patient trajectories. Their method used a name entity recognition (NER) tool to extract medical concepts from discharge summaries, based on which a patient sequence was constructed chronologically. Then they trained a standard GPT model using all patient sequences constructed from the previous step. To forecast future events, they fed a patient history to prompt the model and employed a Monte Carlo sampling strategy to calculate the probability of developing certain conditions. One limitation of this work was that the model could not predict when a certain condition would happen due to a lack of temporal information in their patient sequence. Nevertheless, their work demonstrated the potential of using GPT for modeling patient sequences. We aim to address these limitations in our work.
Figure 1: The OMOP data is first converted to patient sequences by an OMOP encoder based on the patient representation that preserves demographics, visit types, and temporal intervals between visits. Then a generative model is trained to learn the sequence distribution to generate new sequences. Next, the generated sequences are converted back to the OMOP format using an OMOP decoder.
@@ 3 Methods
In Figure 1, we present the framework for generating synthetic EHR data from an OMOP source. To retain the temporal dependencies, we opted to work directly with time-series patient sequences instead of using a bag-of-words (BOW) representation by building on our previous work [14]. We first encoded the OMOP data into patient sequences using a specific patient representation, described later in this section. Secondly, we trained a generative model on the converted patient sequences and utilized it to generate new synthetic patient sequences. Finally, we fed the synthetic patient sequences into an OMOP decoder to create a synthetic OMOP dataset. Furthermore, an evaluation procedure was developed to assess the similarity between the synthetic OMOP and the source OMOP data.
@@ 4.2.1 Dimension-wise Distribution
Figure 3: KL divergence for comparing concept probability distribution between synthetic data and real data. The probabilities of concepts were calculated on the scale of the entire population.
@@
Figure 4: Concept prevalence comparison between the source OMOP and generated OMOP using top $p=95\%$ in the log scale stratified by domain in columns and by population in rows, where x-axis and y-axis represent the source and the synthetic data respectively, and each dot represents a concept
@@ 4.2.2 Co-occurrence Relationship
Figure 5: KL divergence associated with different synthetic data. The closer to the lower bound in the bottom left corner, the better the synthetic data.
Figure 6: Top 100 co-occurring concept pairs for each co-occurrence category e.g. condition-condition interpreted as a condition concept followed by another condition concept. The x and y axes represent the synthetic and source data respectively
@@
Figure 7: Comparison of temporal co-occurrence networks around Type 2 Diabetes Mellitus (T2DM) for real (left) and synthetic data (right). Networks were constructed with T2DM’s (large circle) top 5 co-occurring condition concepts and their own top 5 co-occurring condition concepts. The metrics Prevalence (top) and Pointwise Mutual Information-3 (PMI3) (bottom) were used to quantify co-occurrence. Gold edges indicate edges that are shared by both real and synthetic data, edge thickness indicates the strength of co-occurrence, and arrow direction indicates the direction of the temporal association.
For an in-depth analysis, we focused on examining the co-occurrence relationships associated with Type 2 Diabetes Mellitus (T2DM). Figure 7 qualitatively compares the real and top_p=95synthetic data networks of condition concepts that co-occur around T2DM. The degree of co-occurrence is calculated with prevalence, which evaluates the frequency of concept co-occurrence, and Pointwise Mutual Information-3 (PMI3), which evaluates the probabilistic association of concepts. There is much overlap between the real and synthetic networks when prevalence is used to construct them, but connections are more disparate when PMI3 is used.
@@ 4.2.3 Predictive performance
HF readmission HF patients who have a 30-day all-cause readmission. Observation window: 360 days, Prediction windows 30 days
Hospitalization 2-year risk of hospitalization starting from the 3rd year since the initial entry into the EHR system. Observation window: 540 days, hold-off window: 180 days, prediction windows: 720 days
COPD readmission COPD patients who have a 30-day all-cause readmission. Observation window: 360 days, prediction windows: 30 days
Afib ischemic stroke Afib patients with 1-year risk since the initial diagnosis of afib ischemic stroke. Observation window: 720 days, prediction windows: 360 days
CAD CABG Patients initially diagnosed with Coronary Arterial Disease (CAD) without any prior stent graft will receive the Coronary artery bypass surgery (CABG) treatment. Observation window: 720 days, prediction windows: 360 days
@@
Figure 8: The consolidated distance metrics for different synthetic datasets stratified by cohort. GPT-Vanilla and CEHR-BERT were omitted as the cohorts couldn’t be constructed due to the loss of temporal information. GPT-OUTPAT was omitted from HF readmission and COPD readmission as the cohorts for the same reason.
@@ 4.3 Privacy Evaluations
We adopted the privacy evaluation framework outlined by Yan et al. [10], focusing on quantifying the risk of privacy breaches through various types of attacks, including membership inference, attribute inference, meaningful identity disclosure, and nearest neighborhood adversary attacks (NNAA). We detailed each metric in the subsequent sections and summarized the results in Table 4, which include a comparison to the privacy risk quantification calculated by Yan et al. using the University of Washington (UW) dataset. Our results are comparable to those from all models tested by Yan et al.. Furthermore, according to Yan et al., a risk score below $0.333$ for all defined metrics indicate a low risk to privacy [10]. Note that this evaluation framework requires a separate evaluation dataset not used in training. To ensure a fair evaluation, we held out $11,066$ patients ($0.5\%$ of the training set) to create an evaluation set. Subsequently, we re-trained the model using the remaining patients, following the procedure detailed in Section 4.1. We then generated 1 million synthetic patients using a top_p$=95\%$ specifically for evaluating the privacy metrics.
Privacy Metric CEHR-GPT medGAN medBGAN EMR-WGAN WGAN DPGAN
Attribute Inference 0.027 0.0078 0.0117 0.0680 0.0042 0.0136
Membership Inference 0.1266 0.1506 0.1828 0.2966 0.1758 0.0000
Meaningful Identity Disclosure 0.002 0.0021 0.0027 0.00361 0.0034 0.0004
NNAA Risk -0.0011 0.0008 0.0004 0.0198 0.0085 0.0017
Table 4: Privacy metrics computed on the synthetic data generated by CEHR-GPT. In addition, we include the benchmark metrics for the UW dataset from Yan, C. et al.. [10]
@@ 4.3.1 Membership Inference Attack
In a membership inference attack, the objective is to determine whether a real patient record was utilized in the training of a generative model using the synthetic data [10, 19]. Such attacks can occur if the model overfits to the training data, causing the synthetic data to resemble the training dataset more closely than the non-training dataset.
The attack dataset was constructed as follows: 1) We generated negative examples by randomly selecting 10,000 samples from the holdout set and labeling them as negative; 2) We created positive examples by randomly selecting 10,000 patients from the training set and labeling them as positive. We then matched the patients in the attack dataset to the training dataset by calculating Euclidean distance between the patients in attack dataset and training dataset. Following the guidelines suggested in [10], we used a threshold distance of 5, below which patients were classified as positive (i.e., their data is present in the training dataset) and above which as negative (i.e., their data is not present in the training dataset). This procedure was repeated 100 times to ensure robustness. We evaluated the method by calculating recall, precision, and the F1 score based on the ground truth labels and the inferred predictions. The F1 score was then used to assess the risk associated with the data. The full results can be found in Supplementary Materials 7.4.1.
@@ 4.3.2 Attribute Inference Attack
In the attribute inference attack, we assumed the attacker has access to a group of target patients along with their demographic data, including age, gender, race, the year of their initial clinical visit, and prevalent clinical conditions like hypertension, abdominal pain, chest pain, etc. With these features, the attacker identifies the patient in the synthetic dataset with the closest attribute resemblance to the target individual. Then the attacker aims to use the sensitive attributes of the matched patient from the synthetic dataset to infer the corresponding sensitive attributes of the target patient.
@@
To evaluate the attribute inference risk, we randomly sampled 150,000 patients from the real patient dataset as a target group. The search group comprised 1 million generated synthetic patients. To find the most similar patients from the search group, we created a set of common attributes including the demographic data along with the top 1% of the most prevalent condition concepts, represented as one-hot encoded features. Then a k-nearest neighbors (KNN with k = 1) algorithm was applied to each target patient and a synthetic patient with the smallest Euclidean distance was found. Finally, we extracted the sensitive attributes (condition concepts not in the top 1% tier) from matched target and synthetic patients. F1 scores were computed for the sensitive attributes of each matched patient pair and aggregated across all matched patients.
@@
A baseline analysis was conducted by comparing real patients to other real patients. A result lower than the baseline suggests that the likelihood of finding synthetic data similar to a real patient is lower than finding a real patient similar to a real patient who share sensitive attributes. This implies a lower attribute inference risk which could be acceptable. Our training set was randomly divided into two halves, with 1 million real patients assigned as the target group, and the remaining 1 million as the search group. The matching process was consistently applied and an aggregated F1 score was computed. The full results can be found in Supplementary Materials 7.4.2.
4.3.3 Meaningful Identity Disclosure Attack
The meaningful identity disclosure risk measures the likelihood of identifying a synthetic record by linking it to patient records in an identified population dataset that contains the training data. This linkage could potentially reveal sensitive information about individuals in the population. Adopted from the metrics developed by [20], we can calculate the metrics using the following equation:
$max\Big{(}\frac{1}{N}\sum^{n}_{s=1}(\frac{1}{f_{s}}\times\frac{1+\lambda_{s}}{2}\times I_{s}\times R_{s}),\frac{1}{n}\sum^{n}_{s=1}(\frac{1}{F_{s}}\times\frac{1+\lambda_{s}}{2}\times I_{s}\times R_{s})\Big{)}$
where $N$ denotes the size of the population cohort, $n$ denotes the size of the real sample, $s$ denotes the record index in the real sample, $f_{s}$ represents the number of records matched with record $s$ in the real sample and $F_{s}$ represents the number of records matched with record $s$ in the population cohort. $\lambda_{s}$ is the upper bound of the adjustment parameter for incorrect matches. Following the same configuration in [20], the midpoint between $\lambda_{s}$ and the maximum value of 1 is used to be more conservative. $I_{s}$ is the binary indicator if any real record is matched with the synthetic record. $R_{s}$ is another binary indicator if the adversary could use at least Lsensitive attributes from matched synthetic records to learn something new about the target.
To assess the risk, we implemented the risk model applied by [10] on the CUIMC patient population. We used 3.7M as the population cohort, 2.3M patients as the real sample, and 1M synthetic patients as the target cohort. The matching was performed using 10 QID (Quasi Identifiers) which consist of age, gender, and race demographic attributes and 7 most common diseases in CUIMC patients. Moreover, we established $\lambda_{s}$ at 0.23, setting it as the upper bound to gauge the maximum risk level. Since we have around 17K attributes in total across the condition, drug, and procedure domains, to be aligned with the threshold in [10], we chose 0.1 for L, which resulted in approximately 17 sensitive attributes. We further assessed the risk scores based on the matching across two generalization levels of age group (e.g. age 32 vs age group 30-40) and any combinations of 7 common diseases, resulting in 256 runs. The full results can be found in Supplementary Materials 7.4.3.
4.3.4 Nearest Neighbor Adversary Attack
The Nearest Neighbor Adversary Attack (NNAA) risk is a measure that assesses model overfitting by comparing the closeness of training and synthetic data. If the distance between the training and synthetic data is less than that between the evaluation and synthetic data, it indicates potential overfitting. This concept forms the basis for the metric developed by the researchers. Let’s denote $S_{T}$, $S_{E}$, and $S_{S}$ to be random samples from the training, evaluation, and synthetic datasets. The risk score is defined as the following,
$AA_{ES}-AA_{TS}$
where $AA_{ES}$ and $AA_{TS}$ are defined below,
$AA_{ES}=\frac{1}{2}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}d_{ES}(I)>d_{EE}(i)\big{)}+\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}d_{SE}(i)>d_{SS}(i)\big{)}\Big{)}$
$AA_{TS}=\frac{1}{2}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}d_{TS}(i)>d_{TT}(i)\big{)}+\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{(}d_{ST}(i)>d_{SS}(i)\big{)}\Big{)}$
Here, $d_{ES}(i)=\min_{j}||x_{E}^{i}-x_{S}^{j}||$ indicates the distance between the evaluation data point $x_{E}^{i}\in S_{E}$ and its nearest synthetic neighbor $x_{S}^{j}\in S_{S}$; $d_{EE}(i)=\min_{j,j\neq i}||x_{E}^{i}-x_{E}^{j}||$ represents the distance between the evaluation data point $x_{E}^{i}$ and its nearest neighbor within the same evaluation set $S_{E}$, excluding itself. Analogous definitions apply for $d_{SE}$, $d_{SS}$, $d_{TS}$, $d_{TT}$, $d_{ST}$, and $d_{SS}$. For our study, we randomly chose samples of 10,000 patients from the training, evaluation, and synthetic datasets to compute the metric as per the outlined formula. The full results can be found in Supplementary Materials 7.4.4.